Through what angle will a 5-MeV alpha particle be scattered as it approaches a gold nucleus with an impact parameter of 2.6x10-13m?(Given that Z for Gold=79)

Determine the impact parameter of a 8-MeV alpha particle scattered by 12.5o if it approaches a gold nucleus.
1
Expert's answer
2020-05-05T18:40:50-0400

1)

"b=\frac{kZe^2\cot{0.5\theta}}{K}"

"2.6\cdot 10^{-13}=\frac{(9 \cdot 10^{9})(79)(1.6\cdot 10^{-19})^2\cot{0.5\theta}}{5\cdot 1.6\cdot 10^{-13}}"

"\theta=10\degree"

2)


"b=\frac{(9 \cdot 10^{9})(79)(1.6\cdot 10^{-19})^2\cot{25}}{8\cdot 1.6\cdot 10^{-13}}"

"b=3.05\cdot 10^{-14}m"